# Explainable Ensemble Learning for Graph-Based Malware Detection

### v1

@@ 1 Introduction
Graph Neural Networks (GNNs) have recently emerged as powerful tools for learning from graph-structured data and have demonstrated considerable success in a range of cybersecurity tasks, particularly malware detection [2, 3, 4, 5, 6, 7]. Unlike traditional neural networks, GNNs are specifically designed to operate on graph data by iteratively aggregating information from neighboring nodes and edges, allowing them to capture intricate structural and relational dependencies across the graph. This capability makes GNNs especially well-suited for modeling software behavior, which is often naturally represented as a graph [8]. In malware detection, GNNs have been applied to various graph-based representations of programs, including function call graphs [9], API call graphs [10], and most importantly, control flow graphs (CFGs) [11], each of which captures different behavioral characteristics. Among these representations, CFGs are particularly valuable because they illustrate the execution flow of a program in a structured and analyzable form. A CFG models basic blocks of instructions as nodes and the possible control transfers between them as directed edges. This abstraction captures both high-level logic and low-level execution patterns, enabling the identification of structural anomalies and irregular control paths that are often indicative of malicious behavior.When combined with GNNs, CFGs enable the extraction of rich representations that capture both local execution behaviors and global structural context, thereby improving the detection of sophisticated and evasive malware.
While individual GNN models have shown promising performance, leveraging ensemble learning can further improve generalization, robustness, and predictive performance [12, 13, 14, 15, 16]. Ensemble techniques combine the outputs of multiple base learners to reduce variance and avoid overfitting [17]. Among these methods, stacking stands out as a powerful strategy that trains a meta-learner to combine the predictions of multiple base GNN models. Each GNN in the ensemble may capture different aspects of the input graph due to architectural or training diversity, and the meta-learner can synthesize these diverse outputs into a more accurate and reliable prediction. This hierarchical learning paradigm has been underexplored in malware detection with GNNs and offers a promising direction for improving detection accuracy and resilience against evasive techniques [18, 19, 20, 21, 22].
@@ 2 Related Works
Recent research has increasingly focused on advancing malware detection through graph-based learning techniques, particularly those that model program behavior using CFGs. GNNs have shown strong performance in capturing the structural dependencies within such graphs, enabling robust classification of malicious software. At the same time, ensemble learning approaches, particularly stacking-based frameworks, have been explored to combine the strengths of multiple models for improved robustness and accuracy. As these models grow in complexity, the need for explainability has become more pronounced, especially for deployment in security-critical systems. This section reviews prior work across three major areas: GNN-based malware detection, stacking-based ensemble methods for malware analysis, and explainable GNN frameworks for interpreting malicious behavior.
@@ 3.1 Dynamic CFG Extraction and Node Feature Embedding
CFGs model the execution flow of a program, where nodes represent basic blocks, which are sequential groups of instructions with a single entry and a single exit point. Edges indicate possible transitions in control between these blocks. Static CFGs, constructed through disassembly, often miss important execution paths due to obfuscation, indirect jumps, or dynamic code loading. In contrast, dynamic CFGs are generated by tracing the actual runtime behavior of a program, capturing execution paths that may not be visible through static analysis. This makes them more effective for analyzing advanced or evasive malware. In our framework, we extract dynamic CFGs from PE files to build accurate graph representations for downstream feature embedding and classification.
@@ 3.2 Graph Neural Networks as Base Learners
Using different GNN architectures as base learners enriches the feature space available to the meta-learner. Each GNN type captures distinct structural and semantic patterns due to differences in aggregation functions, update mechanisms, and learning biases. For malware detection, where subtle behavioral variations in control flow graphs can be indicative of malicious activity, such architectural diversity enhances both robustness and classification performance. The meta-learner can leverage these complementary perspectives to improve generalization and decision accuracy.
@@ 3.3 Attention-Based Stacking Ensemble Architecture
Figure 3: SE model training process.
Stacking ensemble (SE) learning is a powerful strategy for improving prediction accuracy and robustness by leveraging the complementary strengths of multiple models. Rather than relying on a single model’s output, stacking combines the predictions of several base learners through a secondary model, known as the meta-learner, to produce a more reliable final decision. This approach enhances generalization and reduces overfitting by mitigating individual model biases and variances. In our work, we adopt an SE framework that integrates multiple GNN-based base learners with an attention-based meta-learner. The training procedure consists of three main phases: (i) independently training each base learner using cross-validation, (ii) generating a new meta-training dataset from the base learners’ validation predictions, and (iii) training the meta-learner using this aggregated information. The overall process is illustrated in Figure 3.
@@
After constructing $\widehat{{Y}}_{i}$ for all $n$ base learners, we horizontally concatenate them to form the input matrix for the meta-learner:
@@ 3.4 Aggregation-Driven Explainability for Stacked GNNs
Post-hoc GNN explainability methods aim to interpret the predictions of trained GNN models by identifying important substructures in the input graph. These methods typically consist of three core components: (i) a pre-trained GNN model whose predictions are to be explained, (ii) a score generation mechanism that produces importance scores over graph elements (nodes, edges, or features), and (iii) an objective function that guides the explainer to identify input elements that most significantly influence the model’s prediction. A variety of explanation techniques have been proposed in the literature, including perturbation-based methods, gradient-based approaches, surrogate modeling techniques, and decomposition-based methods, each offering distinct advantages in terms of fidelity, interpretability, and computational cost. An overview of this process is illustrated in Figure 4.
@@ 4 Results and Analysis
In our experiments, we utilized malicious samples from the BODMAS [33] and PMML [34] datasets, as well as benign samples obtained from the DikeDataset [35]. A summary of the datasets employed in this study is provided in Table 1. For dynamic CFG recovery, we leveraged the Angr framework [36, 37, 38], a Python-based binary analysis tool. Angr facilitates CFG construction by combining symbolic execution with constraint solving, allowing for precise and in-depth graph generation. All experiments were conducted on a workstation equipped with an Intel Xeon Platinum 8253 CPU (32 cores, 3.0 GHz) and 128 GB RAM. We implemented our framework in Python using PyTorch Geometric for model training and NetworkX v2.8.8 for graph processing and manipulation.
@@
Tables 2,3,4, and 5 present the class-wise evaluation metrics, including precision, recall, and F1-score for both benign and malicious classes, along with the overall accuracy of each model. The results cover the performance of the base learners (GCN, GIN, and GAT) as well as the SE model on the test dataset. Furthermore, Figure 6 provides a visual representation of these metrics using a radar plot.
Across all evaluation criteria, including precision, recall, F1-score, and accuracy, the SE model consistently achieves the highest performance. This demonstrates that combining multiple base learners through ensemble learning enhances both generalization and robustness.
@@
Table 2: Test performance of GCN model.
Metric Precision Recall F1-Score Support
Benign 0.7903 0.7656 0.7778 64
Malicious 0.8558 0.8725 0.8641 102
Accuracy 0.8313 166
Macro Avg 0.8230 0.8191 0.8209 166
Weighted Avg 0.8305 0.8313 0.8308 166
Table 3: Test performance of GIN model.
Metric Precision Recall F1-Score Support
Benign 0.7215 0.8906 0.7972 64
Malicious 0.9195 0.7843 0.8466 102
Accuracy 0.8253 166
Macro Avg 0.8205 0.8375 0.8219 166
Weighted Avg 0.8432 0.8253 0.8275 166
Table 4: Test performance of GAT model.
Metric Precision Recall F1-Score Support
Benign 0.8364 0.7188 0.7731 64
Malicious 0.8378 0.9118 0.8732 102
Accuracy 0.8373 166
Macro Avg 0.8371 0.8153 0.8232 166
Weighted Avg 0.8373 0.8373 0.8346 166
Table 5: Test performance of SE model.
Metric Precision Recall F1-Score Support
Benign 0.8475 0.7812 0.8130 64
Malicious 0.8692 0.9118 0.8900 102
Accuracy 0.8614 166
Macro Avg 0.8583 0.8465 0.8515 166
Weighted Avg 0.8608 0.8614 0.8603 166
(a) (a) Benign class performance across models.
(b) (b) Malicious class performance across models.
Figure 6: Class-wise test performance of all models.
@@
The results demonstrate that, for both IG and GBP, our aggregation-based explanation method applied to the SE model consistently achieves lower FidelityFidelity{-} values across different sparsity levels, indicating a stronger ability to identify and prioritize influential subgraphs. Furthermore, in terms of Fidelity+Fidelity{+}, the proposed explanation approach yields values that are comparable to, or slightly higher than, those obtained from the base learners. This suggests that the aggregation method effectively preserves critical information while improving interpretability. Overall, these findings confirm the utility and robustness of the proposed explanation framework in the context of ensemble-based GNN models.

### v2

# Explainable Attention-Guided Stacked Graph Neural Networks for Malware Detection

@@ 1 Introduction
Graph Neural Networks (GNNs) have recently emerged as powerful tools for learning from graph-structured data and have demonstrated considerable success in a range of cybersecurity tasks, particularly malware detection [2, 3]. Unlike traditional neural networks, GNNs are specifically designed to operate on graph data by iteratively aggregating information from neighboring nodes and edges, allowing them to capture intricate structural and relational dependencies across the graph. This capability makes GNNs especially well-suited for modeling software behavior, which is often naturally represented as a graph [4]. In malware detection, GNNs have been applied to various graph-based representations of programs, including function call graphs [5], API call graphs [6], and most importantly, control flow graphs (CFGs) [7], each of which captures different behavioral characteristics. Among these representations, CFGs are particularly valuable because they illustrate the execution flow of a program in a structured and analyzable form. A CFG models basic blocks of instructions as nodes and the possible control transfers between them as directed edges. This abstraction captures both high-level logic and low-level execution patterns, enabling the identification of structural anomalies and irregular control paths that are often indicative of malicious behavior. When combined with GNNs, CFGs enable the extraction of rich representations that capture both local execution behaviors and global structural context, thereby improving the detection of sophisticated and evasive malware.
While individual GNN models have shown promising performance, leveraging ensemble learning can further improve generalization, robustness, and predictive performance [8, 9, 10]. Ensemble techniques combine the outputs of multiple base learners to reduce variance and avoid overfitting [11]. Among these methods, stacking stands out as a powerful strategy that trains a meta-learner to combine the predictions of multiple base GNN models. Each GNN in the ensemble may capture different aspects of the input graph due to architectural or training diversity, and the meta-learner can synthesize these diverse outputs into a more accurate and reliable prediction. This hierarchical learning paradigm has been underexplored in malware detection with GNNs and offers a promising direction for improving detection accuracy and resilience against evasive techniques [12, 13, 14, 15].
@@ 2 Related Works
Recent research has increasingly focused on advancing malware detection through graph-based learning techniques, particularly those that model program behavior using CFGs. GNNs have shown strong performance in capturing the structural dependencies within such graphs, enabling good classification of malicious software. At the same time, ensemble learning approaches, particularly stacking-based frameworks, have been explored to combine the strengths of multiple models for improved robustness and accuracy. As these models grow in complexity, the need for explainability has become more pronounced, especially for deployment in security-critical systems. This section reviews prior work across three major areas: GNN-based malware detection, stacking-based ensemble methods for malware analysis, and explainable GNN frameworks for interpreting malicious behavior.
@@ 3.1 Dynamic CFG Extraction and Node Feature Embedding
CFGs model the execution flow of a program, where nodes represent basic blocks, which are sequential groups of instructions with a single entry and a single exit point. Edges indicate possible transitions between these blocks. Static CFGs, constructed through disassembly, often miss important execution paths due to obfuscation, indirect jumps, or dynamic code loading. In contrast, dynamic CFGs are generated by tracing the actual runtime behavior of a program, capturing execution paths that may not be visible through static analysis. This makes them more effective for analyzing advanced or evasive malware. In our framework, we extract dynamic CFGs from PE files to build accurate graph representations for downstream feature embedding and classification.
@@ 3.2 Graph Neural Networks as Base Learners
Using different GNN architectures as base learners enriches the feature space available to the meta-learner. Each GNN type captures distinct structural and semantic patterns due to differences in aggregation functions, update mechanisms, and learning biases. For malware detection, where subtle behavioral variations in control flow graphs can be indicative of malicious activity, such architectural diversity enhances classification performance. The meta-learner can leverage these complementary perspectives to improve generalization and decision accuracy.
@@ 3.3 Attention-Based Stacking Ensemble Architecture
Figure 3: Training procedure of the attention-based SE model
Algorithm 1 Training the Attention-Based SE model.
1:Input: PE Training Dataset $\mathcal{D}_{\text{train}}$
2:Output: Base learners ($\text{GNN}_{i}$) and meta-learner (attention weights + MLP)
3:function EncodeFeatures(PE)
4:  Extract CFG $G=(V,E)$
5:  $X\leftarrow[\,]$
6:  for $k\in V$ do
7:   Encode instructions to ${e}_{\text{instr}}^{(t)}\in\mathbb{R}^{439}$
8:   ${e}_{\text{node}}^{(k)}\leftarrow\text{AGG}({e}_{\text{instr}}^{(1)},\ldots,{e}_{\text{instr}}^{(n_{k})})$
9:   ${x}^{(k)}\leftarrow f_{\theta}({e}_{\text{node}}^{(k)})\in\mathbb{R}^{d}$
10:   $X\leftarrow[\,X;~x^{(k)}\,]$
11:  end for
12:  return $(G,X)$
13:end function
14:for each $(\text{PE},y)\in\mathcal{D}_{\text{train}}$ do
15:  $(G,X)\leftarrow$ EncodeFeatures(PE)
16:end for
17:Partition $\mathcal{D}_{\text{train}}\to\{\mathcal{D}^{(f)}\}_{f=1}^{5}$
18:for $i=1$ to $n$ do
19:  $Y_{i}\leftarrow[\,]$
20:  for $f=1$ to 55 do
21:   $\text{TrainSet}\leftarrow\mathcal{D}_{\text{train}}\setminus\mathcal{D}^{(f)}$
22:   $\text{ValSet}\leftarrow\mathcal{D}^{(f)}$
23:   Train $\text{GNN}_{i}^{(f)}$ on $(G,X,y)\in\mathcal{D}_{\text{TrainSet}}$
24:   $\widehat{Y}_{i}^{(f)}\leftarrow\text{GNN}_{i}^{(f)}(\text{ValSet})$
25:   $Y_{i}\leftarrow\begin{bmatrix}Y_{i};~\widehat{Y}_{i}^{(f)}\end{bmatrix}$
26:  end for
27:end for
[…]
29:Train meta-learner (attention weights $(w_{i},b_{i})$ and MLP) on $(Y,y)$
30:for $i=1$ to $n$ do
31:  Train $\text{GNN}_{i}$ on $\{(G,X,y)\in\mathcal{D}_{\text{train}}\}$
32:end for
33:return $\text{GNN}_{i}$, $(w_{i},b_{i})$, MLP
Stacking ensemble (SE) learning is a powerful strategy for improving prediction accuracy and robustness by leveraging the complementary strengths of multiple models. Rather than relying on a single model’s output, stacking combines the predictions of several base learners through a secondary model, known as the meta-learner, to produce a more reliable final decision. This approach enhances generalization and reduces overfitting by mitigating individual model biases and variances. In our work, we adopt an SE framework that integrates multiple GNN-based base learners with an attention-based meta-learner. The training procedure consists of three main phases: (i) independently training each base learner using cross-validation, (ii) generating a new meta-training dataset from the base learners’ validation predictions, and (iii) training the meta-learner using this aggregated information. The overall training procedure is illustrated in Figure 3 and summarized in Algorithm 1.
@@
After constructing $\widehat{{Y}}_{i}$ for all $n$ base learners, we concatenate them to form the input matrix for the meta-learner:
@@ 3.4 Aggregation-Driven Explainability for Stacked GNNs
Post-hoc GNN explainability methods aim to interpret the predictions of trained GNN models by identifying important subgraphs in the input graph. These methods typically consist of three core components: (i) a pre-trained GNN model whose predictions are to be explained, (ii) a score generation mechanism that produces importance scores over graph elements (nodes, edges, or features), and (iii) an objective function that guides the explainer to identify input elements that most significantly influence the model’s prediction. A variety of explanation techniques have been proposed in the literature, including perturbation-based methods, gradient-based approaches, surrogate modeling techniques, and decomposition-based methods, each offering distinct advantages in terms of fidelity, interpretability, and computational cost. An overview of this process is illustrated in Figure 4.
@@ 4 Results and Analysis
In our experiments, we utilized malicious samples from the BODMAS [26] and PMML [27] datasets, as well as benign samples obtained from the DikeDataset [28]. A summary of the datasets employed in this study is provided in Table 1. For dynamic CFG recovery, we leveraged the angr framework [29], a Python-based binary analysis tool. Angr facilitates CFG construction by combining symbolic execution with constraint solving, allowing for precise and in-depth graph generation. All experiments were conducted on a workstation equipped with an Intel Xeon Platinum 8253 CPU (32 cores, 3.0 GHz) and 128 GB RAM. We implemented our framework in Python using PyTorch Geometric for model training and NetworkX v2.8.8 for graph processing and manipulation.
@@
Table 2: Test performance of GCN, GIN, GAT, and SE models.
Model Benign Malicious Accuracy
Precision Recall F1 Score Precision Recall F1 Score
GCN 0.7903 0.7656 0.7778 0.8558 0.8725 0.8641 0.8313
GIN 0.7215 0.8906 0.7972 0.9195 0.7843 0.8466 0.8253
GAT 0.8364 0.7188 0.7731 0.8378 0.9118 0.8732 0.8373
SE 0.8475 0.7812 0.8130 0.8692 0.9118 0.8900 0.8614
Table 2 presents the class-wise evaluation metrics, including precision, recall, and F1-score for both benign and malicious classes, along with the overall accuracy of each model. The results cover the performance of the base learners (GCN, GIN, and GAT) as well as the SE model on the test dataset. Furthermore, Figure 6 provides a visual representation of these metrics using a radar plot.
(a) Benign class performance across models.
(b) Malicious class performance across models.
Figure 6: Class-wise test performance of all models.
Across all evaluation criteria, including precision, recall, F1-score, and accuracy, the SE model consistently achieves the highest performance. This demonstrates that combining multiple base learners through ensemble learning enhances generalization.
@@
The results demonstrate that, for both IG and GBP, our aggregation-based explanation method applied to the SE model consistently achieves lower FidelityFidelity{-} values across different sparsity levels, indicating a stronger ability to identify and prioritize influential subgraphs. Furthermore, in terms of Fidelity+Fidelity{+}, the proposed explanation approach yields values that are comparable to, or slightly higher than, those obtained from the base learners. This suggests that the aggregation method effectively preserves critical information while improving interpretability. Overall, these findings confirm the utility of the proposed explanation framework in the context of ensemble-based GNN models.